MILLIMETER DETECTION OF AlO (X^{2}Σ^{+}): METAL OXIDE CHEMISTRY IN THE ENVELOPE OF VY CANIS MAJORIS

E. D. TENENBAUM, \underline{L. M. ZIURYS}, University of Arizona, Steward Observatory, Department of Chemistry, Arizona Radio Observatory Tucson, AZ 85721.

A new circumstellar molecule, the radical AlO (X^{2}Σ^{+}), has been detected toward the envelope of the oxygen-rich supergiant star VY Canis Majoris (VY CMa) using the Arizona Radio Observatory (ARO). The N = 7 → 6 and 6 → 5 rotational transitions of AlO at 268 and 230 GHz were observed at 1 mm using the ARO Submillimeter Telescope (SMT) and the N = 4 → 3 line was detected at 2 mm using the ARO 12 m. Based on the shape of the line profiles, AlO most likely arises from the dust-forming region in the spherical outflow of VY CMa, as opposed to the blue- or red-shifted winds, with a source size of θ_{s} ∼ 0.5′′. Given this source size, the column density of AlO was found to be N_{tot} ∼ 2 × 10^{15} cm^{−2} for T_{rot} ∼ 230 K, with a fractional abundance, relative to H_{2}, of ∼ 10^{−8}. Gas-phase thermodynamic equilibrium chemistry is the likely formation mechanism for AlO in VY CMa, but shocks may disrupt the condensation process into Al_{2}O_{3}, allowing AlO to survive to a radius of ∼ 20 R_{∗}. The detection of AlO in VY CMa is additional evidence of an active gas-phase refractory chemistry in oxygen-rich envelopes, and suggests such objects may be fruitful sources for other new oxide identifications.